OpenStudy (anonymous):

a solution is made by dissolving 34 g of NH3 in 2.00 mL of water. calculate the molality of the solution. answer is 1.02 but i got 0.998. pls show me how to calculate this the right way :/

OpenStudy (anonymous):

n(NH3)= m/M =34/(3x1.008+14.01) =1.996 mol molality=1.996/2 =0.998

OpenStudy (anonymous):

yeah i got that to. so am i correct? or wrong? or are the answer is just wrong?

OpenStudy (anonymous):

yea you should be right~ :)

OpenStudy (anonymous):

thnks!